# GFlowGR: Fine-tuning Generative Recommendation Frameworks with Generative Flow Networks

### v1

@@ 3.1. Adaptive Trajectory Sampler
The sparsity of recommendation data is a fundamental challenge, as only a minimal fraction of user-item interactions are observed (Gao et al., 2022). Consequently, constructing training trajectories solely from collected data is insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset. A naive approach to address this is through data augmentation (Dang et al., 2024) or random negative sampling (Ma et al., 2024)—for instance, randomly selecting $K$ augmented samples $\{v^{\prime}_{k}\}_{k\leq K}$ such that $v\neq v^{\prime}_{k}$ for all $k$ (Yang et al., 2020; Wang et al., 2021b). However, such randomness introduces training instability and often yields samples misaligned with the model’s evolving learning dynamics. For example, an obviously irrelevant augmented sample that was effective as a negative in early training epochs may become uninformative in later stages, as the model learns to easily distinguish it from true positives.
@@
To address this, we propose an adaptive trajectory sampler that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner. We assume that during the early training phase, LLMs struggle to distinguish between positive and augmented samples, while in later stages, overly simple samples contribute minimally to model improvement (Wang et al., 2021a). Inspired by recent work demonstrating the benefits of integrating collaborative knowledge into LLM-based recommendation systems (Lin et al., 2025a), we leverage a pretrained collaborative model (CM) to guide sample augmentation. Specifically, we score $N$ candidate augmentations $\{v^{\prime}_{n}\}_{n\leq N}$ using $\hat{y}_{CM}^{v^{\prime}_{n}}=\mathrm{CM}(v^{\prime}_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v^{\prime}_{n}}$ represents the collaborative score of item $v^{\prime}_{n}$ to user $U$.
@@
During initial training epochs, we prioritize sampling items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire. As training progresses, the sampler gradually incorporates harder augmented samples with higher collaborative scores—for example, the blue jacket in the visualized example, which aligns more closely with the user’s preferences but remains unobserved in historical data. This curriculum-based approach enables the sampler to adaptively select trajectories rooted in collaborative knowledge, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
3.2. Comprehensive Reward Model
@@
To address this, we introduce a comprehensive reward model that integrates three complementary signals: augmentation status, collaborative scores, and token similarity. For simplicity, we describe the framework using a single augmented sample ($K=1$, easily extendable to multiple samples). Let $\tau=\{s_{l}\}_{l\leq L}$ denote the trajectory of the observed positive item $v$, and $\tau^{\prime}=\{s^{\prime}_{l}\}_{l\leq L}$ denote the trajectory of an augmented item $v^{\prime}$. Three aspects of reward are defined as:
  • Augmentation Signal ($r_{a}$): marks whether a sample is observed (positive) or augmented, i.e., $r_{a}=1$ for the ground-truth trajectory $\tau$ and $r^{\prime}_{a}=0$ for the augmented trajectory $\tau^{\prime}$.
  • Collaborative Score ($r_{c}$): the collaborative score from the same model, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$.
  • Token Similarity ($r_{\text{sim}}$): rewards partial correctness by measuring the number of shared tokens with the positive sample. Specifically, for $\tau^{\prime}$, $r^{\prime}_{\text{sim}}=\sum_{l=1}^{L}\mathbb{I}(t_{l}=t^{\prime}_{l})$, where $\mathbb{I}(\cdot)$ is an indicator function. And $r_{\text{sim}}=L$ for $\tau$. For the visualized example, $r^{\prime}_{\text{sim}}=1$, as it has one shared token with the ground truth.
We sum up these reward signals as the final reward for each trajectory, which can be formally expressed as $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$.
3.3. Training and Inference
Training. The training objective of GFlowGR integrates two complementary loss components: the vanilla next-token prediction loss $\mathcal{L}_{\text{GR}}$ for positive samples, the GFlowNet training loss $\mathcal{L}_{\text{GFN}}$ over sampled trajectories. The inclusion of $\mathcal{L}_{\text{GR}}$ (Equation (1)) ensures the model retains foundational recommendation capabilities while adapting the LLM to GR tasks. For $\mathcal{L}_{\text{GFN}}$, substituting the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ and backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ into Equation (3) and (4) yields:
@@ 3.3. Training and Inference
$\mathcal{F}$, and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5). $\mathcal{L}_{\text{GFN}}$ can be instantiated using either $\mathcal{L}_{\text{DB}}$ or $\mathcal{L}_{\text{TB}}$. The complete loss function for $(U,v,\tau)$ is:
(7) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\mathcal{L}_{\text{GFN}}(\kappa)$
where $\kappa$ represents either the positive trajectory $\tau$ or an augmented trajectory $\tau^{\prime}$, and $\lambda$ is the hyperparameter to balance different objectives. To mitigate exposure bias, we weight each trajectory by the inverse disagreement between the augmentation signal $r_{a}$ and the collaborative score $r_{c}$:
(8) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\frac{1}{(r_{a}-r_{c})^{2}}\mathcal{L}_{\text{GFN}}(\kappa)$
This heuristic weighting strategy downweights examples where the collaborative model exhibits high uncertainty (large $|r_{a}-r_{c}|$), which often correspond to hard negatives or ambiguous interactions (Chen et al., 2023).
@@
Model Finetuning Beauty Yelp
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0204 0.0334 0.0132 0.0174
PPO 0.0020 0.0032 0.0013 0.0017 0.0004 0.0005 0.0003 0.0003
GRPO 0.0347 0.0543 0.0229 0.0291 0.0193 0.0328 0.0126 0.0169
DPO 0.0308 0.0483 0.0200 0.0256 0.0100 0.0182 0.0060 0.0087
S-DPO 0.0314 0.0492 0.0208 0.0265 0.0212 0.0333 0.0137 0.0176
SPRec 0.0312 0.0487 0.0205 0.0258 0.0103 0.0186 0.0067 0.0090
IPA 0.0306 0.0481 0.0198 0.0245 0.0101 0.0179 0.0066 0.0093
GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0221 0.0345 0.0145 0.0185
GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0256* 0.0404* 0.0166* 0.0213*
LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0205 0.0341 0.0131 0.0175
PPO 0.0128 0.0206 0.0087 0.0112 0.0002 0.0004 0.0001 0.0002
GRPO 0.0384 0.0592 0.0251 0.0317 0.0228 0.0380 0.0146 0.0195
DPO 0.0358 0.0571 0.0238 0.0307 0.0172 0.0286 0.0115 0.0151
S-DPO 0.0390 0.0607 0.0259 0.0336 0.0247 0.0389 0.0161 0.0206
SPRec 0.0370 0.0573 0.0249 0.0315 0.0190 0.0303 0.0124 0.0160
IPA 0.0367 0.0586 0.0252 0.0320 0.0197 0.0293 0.0129 0.0163
GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0252 0.0399 0.0166 0.0213
GFlowGR-TB 0.0417* 0.0632* 0.0276* 0.0345* 0.0261* 0.0413* 0.0173* 0.0222*
Inference. The inference process remains unchanged from standard GR frameworks: user prompts are constructed as before, and the fine-tuned LLM generates item identifiers sequentially. By design, GFlowGR enhances GR without altering the inference pipeline, ensuring seamless integration into existing systems.
@@ 4.1. Experimental Setups
Datasets. We used two real-world datasets to evaluate GFlowGR, namely Yelp and Amazon Beauty. The data processing and split are following previous GR works (Wang et al., 2024a; Zheng et al., 2024).
Baselines and Backbones. We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: PPO (Schulman et al., 2017), GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n. d.]), SPRec (Gao et al., 2025a), IPA (Deng et al., 2025). Notably, the specially designed auxiliary loss in addition to SFT is regarded as a part of the GR backbones. And we install GFlowGR to two backbones: TIGER (Rajput et al., 2023) and LETTER (Wang et al., 2024a).

### v2

@@ 3.1. Trajectory Sampler
2. Random sampling. The sparsity of recommendation data is a fundamental challenge, as only a minimal fraction of user-item interactions are observed (Gao et al., 2022). Consequently, constructing training trajectories solely from collected data is sometimes insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset. Consequently, we can also conduct random negative sampling (Ma et al., 2024) for the trajectory sampler. For instance, randomly selecting N1N-1 augmented samples $\{v_{n}\}_{2\leq n\leq N}$ from the item set such that $v_{m}\neq v_{n}$ for all $1\leq m,n\leq N,m\neq n$ (Yang et al., 2020; Wang et al., 2021b).
However, such randomness introduces training instability and often yields samples misaligned with the model’s evolving learning dynamics. For example, an obviously irrelevant augmented sample that was effective as a negative in early training epochs may become uninformative in later stages, as the model learns to easily distinguish it from true positives.
@@
To address this, we propose two adaptive trajectory samplers that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner. We assume that during the early training phase, LLMs struggle to distinguish between positive and augmented samples, while in later stages, overly simple samples contribute minimally to model improvement (Wang et al., 2021a).
@@
3. CM-based sampling. Inspired by recent work demonstrating the benefits of integrating collaborative knowledge into LLM-based recommendation systems (Lin et al., 2025a), we leverage a pretrained collaborative model (CM) to guide sample augmentation. Specifically, we score 2N12N-1 candidate augmentations $\{v_{n}\}_{2\leq n\leq 2N}$ using $\hat{y}_{CM}^{v_{n}}=\mathrm{CM}(v_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v_{n}}$ represents the collaborative score of item $v_{n}$ to user $U$. During initial training epochs, we prioritize sampling N1N-1 items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire. As training progresses, the sampler gradually incorporates harder augmented samples with higher collaborative scores—for example, the blue jacket in the visualized example, which aligns more closely with the user’s preferences but remains unobserved in historical data.
4. On-policy sampling. An instant alternative to CM-based sampling is the on-policy sampling (Deng et al., 2025; Gao et al., 2025a). Specifically, the trajectory sampler can feed the user prompt to current LLM to generate top-(N1)(N-1) negatives for each $U$ with beam search.
These curriculum-based approach enables the sampler to adaptively select trajectories, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
Algorithm 1 Train and inference process of GFlowGR with CM.
1:Indicate the tokenizer.
2:Indicate the LLMs.
3:Indicate the collaborative model ($\mathrm{CM}$).
Train Process
5:Train the tokenzier with $\mathcal{L}_{T}$ and represent items as identifiers with the trained tokenizer.
6:Indicate the hyperparameters for the training, including the number of augmented trajectories (NN), the weight of GFlowNets loss ($\lambda$).
7:for a batch of samples $(U,v)$ do
8:  Construct the user prompt with item tokenizer
9:  Get the trajectory $\tau_{1}$ for the ground-truth $v$.
10:  Generate N1N-1 augmented trajectories $\{\tau_{n}\}_{2\leq n\leq N}$ with the trajectory sampler with one of the strategy described in Section 3.1.
11:  Calculate the step-wise forward probability $\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ of LLM for all trajectories.
12:  Calculate the step-wise flow $\mathcal{F}(s_{l})$ with the flow estimator.
13:  Calculate the reward $\mathcal{R}(s_{L})$ for the terminal state.
14:  Calculate $\mathcal{L}_{\text{GFN}}$ with Equation (5) or Equation (6).
15:  Calculate the SFT loss $\mathcal{L}_{\text{GR}}$ based on Equation (1).
16:  Sum the SFT loss and GFlowNets loss for the final loss $\mathcal{L}$ as in Equation (7).
17:  Update LLM parameters.
18:end for
Inference Process
14:Load the LLM checkpoint and item tokenizer.
15:for arbitary user $\mathcal{U}$ do
16:  Construct the user prompt with item tokenizer.
17:  Generate $L$ tokens as the identifier of the next item.
18:  Decode the identifier and return the predicted item.
19:end for
3.2. Reward Model
@@
To address this, we introduce a comprehensive reward model that can integrate multiple reward signals. For simplicity, we describe the framework using a single augmented sample ($N=2$, easily extendable to multiple samples). Let $\tau=\{s_{l}\}_{l\leq L}$ denote the trajectory of the observed positive item $v$, and $\tau^{\prime}=\{s^{\prime}_{l}\}_{l\leq L}$ denote the trajectory of an augmented item $v^{\prime}$. Several aspects of reward might be defined as:
  • Interaction Signal ($r_{a}$): marks the interaction level for items. For example, $r_{a}=10$ for the liked item trajectory $\tau$, $r_{a}=1$ for the clicked item trajectory $\tau$, and $r^{\prime}_{a}=0$ for the augmented or unpresented trajectory $\tau^{\prime}$.
  • Estimated Score ($r_{c}$): the collaborative score from CM, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$ or the LLM score score (forward probability).
  • Business Target: customized reward for differnet bussiness targets (Zhou et al., 2025). We illustrate with the format reward ($r_{\text{sim}}$), which reveal partial correctness by measuring the number of shared tokens with the positive sample. Specifically, for $\tau^{\prime}$, $r^{\prime}_{\text{sim}}=\sum_{l=1}^{L}\mathbb{I}(t_{l}=t^{\prime}_{l})$, where $\mathbb{I}(\cdot)$ is an indicator function. And $r_{\text{sim}}=L$ for $\tau$. For the visualized example, $r^{\prime}_{\text{sim}}=1$, as it has one shared token with the ground truth.
3.3. GFN Training Objective
The training objective of GFlowGR integrates two complementary loss components: the vanilla next-token prediction loss $\mathcal{L}_{\text{GR}}$ for positive samples, the GFlowNet training loss $\mathcal{L}_{\text{GFN}}$ over sampled trajectories. The inclusion of $\mathcal{L}_{\text{GR}}$ (Equation (1)) ensures the model retains foundational recommendation capabilities while adapting the LLM to GR tasks. For $\mathcal{L}_{\text{GFN}}$, substituting the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ and backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ into Equation (3) and (4) yields:
@@ 3.3. GFN Training Objective
$\mathcal{F}$ and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5). $\mathcal{L}_{\text{GFN}}$ can be instantiated using either $\mathcal{L}_{\text{DB}}$ or $\mathcal{L}_{\text{TB}}$. The complete loss function for $(U,v,\{\tau_{n}\}_{n\leq N})$ is:
(7) $\mathcal{L}(U,v,\{\tau_{n}\}_{n\leq N})=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{n=1}^{N}\mathcal{L}_{\text{GFN}}(\tau_{n})$
where $\lambda$ is the hyperparameter to balance different objectives.
3.4. Training and Inference Process
We delineate the training and inference procedures of GFlowGR in Algorithm 1, with the use of CM. Notably, GFlowGR can still work without CM. Initially, the employed item tokenizer, generative LLMs for GR frameworks, and the collaborative model for GFlowGR are declared (lines 1 - 3). Subsequently, we train the item tokenizer and utilize the trained tokenizer to transform items into identifiers consisting of $L$ tokens (line 5). Prior to fine-tuning LLMs with GFlowGR, we declare the essential hyperparameters (line 6). For a batch of training samples, we first structure the LLM input (line 8) and convert the item $v$ into the generation trajectory $\tau$ (line 9). Next, GFlowGR generates several augmented trajectories (line 10) and computes GFlowNets elements, which include the forward probability, flow, and reward (lines 11 - 13). Finally, the LLM parameters are updated by the sum of the SFT loss and the GFlowNets loss (lines 14 - 17). During the inference phase, we structure the LLM input, generate the subsequent tokens with the checkpoint, and decode the outcome for recommendation (lines 16 - 18), where GFlowGR is not involved.
Table 1. The statistics of the preprocessed datasets
Dataset # Users # Items Sparsity Avg.length
Yelp 30,431 20,032 99.85% 9.39
Beauty 22,362 12,101 99.92% 7.87
Instruments 24,772 9,922 99.99% 7.32
@@ 3.4. Training and Inference Process
Model Finetuning Beauty Instruments Yelp
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0864 0.1071 0.0740 0.0806 0.0204 0.0334 0.0132 0.0174
GRPO 0.0347 0.0543 0.0229 0.0291 0.0834 0.0988 0.0718 0.0768 0.0193 0.0328 0.0126 0.0169
DPO 0.0308 0.0483 0.0200 0.0256 0.0803 0.0977 0.0692 0.0753 0.0100 0.0182 0.0060 0.0087
S-DPO 0.0314 0.0492 0.0208 0.0265 0.0791 0.0930 0.0688 0.0733 0.0212 0.0333 0.0137 0.0176
SPRec 0.0312 0.0487 0.0205 0.0258 0.0773 0.0903 0.0662 0.0709 0.0103 0.0186 0.0067 0.0090
IPA 0.0306 0.0481 0.0198 0.0245 0.0769 0.0899 0.0661 0.0707 0.0101 0.0179 0.0066 0.0093
GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0868 0.1074 0.0747 0.0807 0.0221 0.0345 0.0145 0.0185
GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0882* 0.1093* 0.0747* 0.0815* 0.0256* 0.0404* 0.0166* 0.0213*
LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0853 0.1049 0.0734 0.0797 0.0205 0.0341 0.0131 0.0175
GRPO 0.0384 0.0592 0.0251 0.0317 0.0819 0.0976 0.0709 0.0759 0.0228 0.0380 0.0146 0.0195
DPO 0.0358 0.0571 0.0238 0.0307 0.0788 0.0960 0.0682 0.0737 0.0172 0.0286 0.0115 0.0151
S-DPO 0.0390 0.0607 0.0259 0.0336 0.0769 0.0914 0.0674 0.0722 0.0247 0.0389 0.0161 0.0206
SPRec 0.0370 0.0573 0.0249 0.0315 0.0754 0.0896 0.0655 0.0701 0.0190 0.0303 0.0124 0.0160
IPA 0.0367 0.0586 0.0252 0.0320 0.0750 0.0891 0.0648 0.0699 0.0197 0.0293 0.0129 0.0163
GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0855 0.1071 0.0738 0.0802 0.0252 0.0399 0.0166 0.0213
GFlowGR-TB 0.0433* 0.0672* 0.0286* 0.0363* 0.0888* 0.1100* 0.0750* 0.0819* 0.0261* 0.0413* 0.0173* 0.0222*
@@ 4.1. Experimental Setups
Datasets. In this paper, we conduct comprehensive experiments on three widely-used datasets, i.e., Yelp, Beauty, and Instruments. Yelp [^2] https://www.yelp.com/dataset records user check-ins at restaurants, along with corresponding reviews and ratings. Beauty and Instruments are subcategories from Amazon Review [^3] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews (McAuley et al., 2015), which include user histories on Amazon products. For data preprocessing, we follow the procedures in LETTER (Wang et al., 2024a). Items are tokenized based on titles and descriptions for Beauty and Instruments, and based on titles, descriptions, brands, and categories for Yelp. For data splitting, we take the last interacted item of each user as the test sample, the second last one as the validation, and iterate the remaining part as training samples, i.e., each interacted data is the target for the preceding user sequence. We present the data statistics in Table 1.
Baselines and Backbones. We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n. d.]), SPRec (Gao et al., 2025a), IPA (Deng et al., 2025). Notably, the specially designed auxiliary loss in addition to SFT is regarded as a part of the GR backbones. And we install GFlowGR to two backbones: TIGER (Rajput et al., 2023) and LETTER (Wang et al., 2024a). Detailed introduction is as follows:
  • SFT. Supervised fine-tuning, optimizes LLMs with the next token prediction loss as in Equation (1) of the main text.
  • GRPO (Shao et al., 2024). Group Relative Policy Optimization is another on-policy RL algorithm based on the idea of PPO. GRPO improves the PPO by updating the model with the relative advantages of a group of actions, reducing the dependency on the absolute reward function.
  • DPO (Rafailov et al., 2023). Direct Policy Optimizaion is an off-policy RL algorithm, which updates the model with off-policy pairwise preference data to simplify training while achieving competitive results in alignment tasks.
  • S-DPO (Chen et al., [n. d.]). This is a DPO-like algorithm tailored for language model-based recommendation scenarios. Unlike standard DPO, it can utilize multiple negative samples, making it better adapted to the characteristics of recommendation tasks.
  • SPRec (Gao et al., 2025a). This method selects rejected samples for DPO to mitigate item-level and token-level biases during LLM fine-tuning for recommendation tasks. Specifically, it updates rejected samples of the off-policy data as negative samples with the highest beam search scores and fine-tunes LLMs via iterative SFT and DPO.
  • IPA (Deng et al., 2025). This method is part of OneRec, which aims to align LLMs with a reward model integrating collaborative knowledge. Similar to SPRec, IPA iteratively updates the model via SFT and DPO steps and replaces rejected samples with those having the highest reward at each step.
@@
Implementation Details. We conduct all experiments on the same GPU device. In terms of the hyperparameter settings, we find the optimal setting for both baselines and GFlowGR. The default ensemple of GFlowGR takes the CM-based sampling for trahectory sampler, and the sum of three reward signals as the reward, i.e., $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$. The training procedure is controlled by Transformer Trainers with early stopping. The implementation of baselines is based on TRL (GRPO, DPO) and the released code of papers (S-DPO, SPRec). We implement IPA by ourselves, where the reward model is the same collaborative model used for GFlowGR.
Figure 3. Parameter study on NN and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to NN.
Table 3. Ablation study with TIGER on Beauty.
Model Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj GFlowGR
TIGER R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0407
R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0651
N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0266
N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0344
LETTER R@5 0.0387 0.0408 0.0384 0.0398 0.0404 0.0410 0.0424 0.0423 0.0433
R@10 0.0596 0.0631 0.0580 0.0610 0.0603 0.0622 0.0659 0.0639 0.0672
N@5 0.0254 0.0265 0.0246 0.0259 0.0263 0.0274 0.0278 0.0274 0.0286
N@10 0.0322 0.0337 0.0310 0.0327 0.0327 0.0342 0.0354 0.0343 0.0363